3[3(X+5)+8X+7]+5[3(X-6)-2(4X-3)]<3X+19=
Answer by Boreal(15235)

3[3(X+5)+8X+7]+5[3(X-6)-2(4X-3)]<3X+19=
3(3x+15+8x+7] + 5{3x-18-8x+6)<3x+19
9x+45+24x+21+15x-90-40x+30<3x+19
8x+6<3x+19
5x<13
x<(13/5)
2 should work
3(21+23) + 5(-12-10)=132-110=22<25